A wave is modeled with the function y(x,t) = (0.66m) cos((0.03rad/m)x − (0.9rad/s)t + π/3). Find the wavelength.




Round your answer to 2 decimal places.



Do not forget the unit

1
Expert's answer
2022-05-12T15:09:26-0400
"y(x,t) = (0.66m)\cos((0.03rad/m)x\\ − (0.9rad/s)t + π/3)"

"y(x,t) = y_0\cos((2\pi/\lambda)x −\omega t + \phi_0)"

Hence, the wavelenght

"\lambda=\frac{2\pi}{0.03}=209.33\:\rm m"